OpenStudy (anonymous):

\[\log_{6}(a ^{2}+2)+\log_{6}2=2\]

OpenStudy (jdoe0001):

$$ log_6(a^2+2)+log_6(2)=2\\ \implies log_6\pmatrix{(a^2+2)\times 2} =2 \implies log_6(2a^2+4)=2\\ \implies 6^{log_6(2a^2+4)}=6^2\\ \text{using the log cancellation rules}\\ \implies 2a^2+4 = 6^2 $$

OpenStudy (jdoe0001):

and I think you can get it from there

OpenStudy (whpalmer4):

Adding logarithms is the same as taking the logarithm of the product. You can rewrite that equation as \[\log_6(2(a^2+2)) = 2\]Then raise 6 to the power of each side, and remember that \(b^{\log_b a} = a\) so you have\[2(a^2+2) = 6^2\]

OpenStudy (anonymous):

x=4

OpenStudy (jdoe0001):

yes

OpenStudy (whpalmer4):

and...

OpenStudy (anonymous):

x=-4

OpenStudy (whpalmer4):

right.